# Dialetto, ma Quanto Dialetto? Transcribing and Evaluating Dialects on a Continuum

### v1

@@ 3.4.4 Ordinary Kriging
Kriging uses the variogram to calculate weights that account for both distance and spatial correlation, providing more accurate estimates at unsampled locations. In ordinary kriging, the value at an unknown location $x_{0}$, denoted as $\hat{Z}(x_{0})$, is a weighted sum of the known values $Z(x_{i})$ at nearby locations:
@@ 3.4.5 Regression Kriging
Regression kriging extends ordinary kriging by incorporating auxiliary variables to account for trends in the data that might otherwise violate the stationarity assumption. These auxiliary variables are commonly assumed to have a linear relationship with the variable of interest.
@@ 4.1 Speech-to-Text Evaluation
[…]
@@
[…]
@@
Figure 3: Effect of training data size on geostatistical interpolation. Orange is nearest neighbor interpolation; black is inverse distance weighting; blue is regression kriging.
The chrF results are geographically interpolated and plotted in Figure 2. We observe in the map that zero-shot performance is particularly high in the Tuscan region, where the highest performing dialect site also resides, as shown in Table 5 and Table 6. Prior literature has established standard Italian to be modelled after Tuscan varieties (Hall, 1980; Wieling et al., 2014), which corroborates the trend observed in our results. The similarity of Umbrian dialects with Tuscan dialects is also observable in Umbrian dialects ranking second after Tuscan in Table 5 and Table 6. Table 3 details the correlation between linguistic similarity and the performance scores, where the Pearson correlation for the chrF score is -0.50 and for BLEU -0.54, and the Spearman correlation is -0.58 for chrF and -0.50 for BLEU, suggesting a strong correlation between similarity to the standard (as approximated by the highest performing site) and speech-to-text performance.
@@ 4.3 Geostatistical interpolation
Building on the observation that there is a clear geographical signal in both the performance and dialectometry maps, we next measure to what extent the incorporation of geographical knowledge helps predict zero-shot speech-to-text performance at unseen sites. As shown in Table 4, geostatistical interpolation is highly predictive of both BLEU and chrF scores, with RMSE scores going to as low as 4.19 and 5.70 by regression kriging, our best method. Figure 3 additionally show the effect of training data size on the regression prediction performance. In all results, we observe that the incorporation of distance and covariance between samples as weighting improves over an interpolation by only the value of the nearest neighbor, where regression kriging is consistently the best performing method, followed by inverse distance weighting.
@@ 5.1 Evaluating Dialects as a Continuum
In Figure 2, we observe that for Italian dialects, the evaluation results stand in a continuum that bears similarity to the map of dialect similarity relations. Furthermore, even when one restricts the examined samples to Tuscan varieties, we observe in Table 5 and Table 6 that there is still a disparity that extends the further away one is from the highest performing sites. Our results highlight the need for work in dialect evaluation to take into account the continuous nature of dialects that may exhibit even within dialects that might fall under the same category. While prior work has made important headway in individually evaluating non-standard varieties as as separate categories and highlighting the performance disparity when compared against the standard, regional variation also exhibits within non-standard varieties such as AAVE (Jones, 2015). Our results suggest that performance prediction in AAVE would arguably also exhibit regional differences, potentially patterning based on how similar the regional varieties are to standard English (e.g. in urban sites), although empirical work is needed to confirm this hypothesis. Importantly, our findings highlight that an evaluation of dialects that is insufficiently balanced geographically therefore carries the risk of overly optimistic views towards model performance at geographically marginal sites, which in turn may lead to social harm towards the subgroups which speak it.
@@ 5.2 Geo-Based Performance Prediction
Extending on the claim of Ahuja et al. (2022), who highlight the role of geography in predicting how well the pivot language for finetuning generalizes to the target language, we explicitly utilize geostatistical methods in our work on geographically proximate dialects, and find both the distance between dialect sites and the covariance between sites to be useful for predicting zero-shot ASR performance at unseen sites. Our results emphasize the geographical structure of dialects, and points to the possibility of leveraging such geographical structure for multilingual transfer between dialects.
@@ 6 Related Work
Kantharuban et al. (2023) stands as the work most similar to our own, where LLMs for both speech and for text are used to evaluate on the tasks of Machine Translation and Automatic Speech Recognition on regional dialects of high and low-resource languages. They find model performance on such varieties to be highly correlated with lexical and phonetic similarity to the highest performing variety. Their work however considers dialect varieties categorically by considering regional varieties such as Argentinian and Chilean Spanish under the same language of Spanish. Our work instead proposes to work on varieties on a continuous basis across geographically nearby varieties.
As regards treating varieties as a continuum, in NLP, Grieve et al. (2024) hypothesize that language models inherently model varieties of language, and propose to leverage sociolinguistic expertise to identify underrepresented varieties. In dialectology the notion of dialects as a continuum is well-established, where computational work in quantifying such a continuum abounds in the field of dialectometry (Heeringa and Nerbonne, 2001; Wieling and Nerbonne, 2015; Nerbonne, 2010).
With regard to bias against dialect varieties in ASR, Feng et al. (2021) show that Dutch ASR systems perform worse on Flanders speakers compared to speakers of all regions in the Netherlands. Kulkarni et al. (2024) show different ASR systems to exhibit different performance biases across 11 states in Brazil, where their dataset consists of scripted speech, and with categorical boundaries between states. Chang et al. (2024) document how self-supervised representations still exhibit a performance disparity in ASR upon AAVE. Our work shows that there is geographical structure in dialect ASR bias that is continuous and correlated with social variables, enabling more fine-grained studies on ASR bias in other languages.
@@ 7 Conclusion
In this paper, we conceptualize dialect relations as a continuum. We cross-examine our results against established research in linguistic dialectometry (Wieling and Nerbonne, 2015). We find zero-shot performance of ASR systems on dialects to pattern similarly to a measure of similarity to the standard variety, observing strong correlations. Furthermore, we introduce geostatistical methods that are predictive of zero-shot performance at held-out sites. Our work highlights the need for more research on non-standard varieties that takes into account the continuum nature of dialects. Doing so holds the potential for uncovering bias and closing prior evaluation gaps.
@@ 8 Limitations
Our study focuses on related varieties of Italian dialects, which potentially limits the generalizability of our findings. The scope of language varieties examined in our study is constrained by the scarcity of geo-tagged dialect datasets, which potentially limits the generalizability of our results to other languages. Furthermore, we approximate the standard variety by the best performing dialect variety, which may affect the correlation results depending on how similar the best performing Tuscan variety actually is to standard Italian. Future work should examine how well our insights generalize to other dialect continuua and to other modalities, where a continuum-based disparity may likewise also exhibit in text-based regional linguistic data.
@@ 9 Acknowledgements
We would like to thank the members of the MaiNLP lab and the anonymous reviewers for their invaluable feedback. The icons featured in Figure 2 are from Freepik (Italy, neural network) [^6] https://www.flaticon.com and SVG Repo (pin, arrow, sound wave) [^7] https://www.svgrepo.com. Lastly, we recognize the support for SS and BP through the ERC Consolidator Grant 101043235.

### v2

@@ 3.3.4 Ordinary Kriging
Kriging (Oliver and Webster, 1990) uses the variogram to calculate weights that account for both distance and spatial correlation, providing more accurate estimates at unsampled locations. In ordinary kriging, the value at an unknown location $x_{0}$, denoted as $\hat{Z}(x_{0})$, is a weighted sum of the known values $Z(x_{i})$ at nearby locations:
@@ 3.3.5 Regression Kriging
Regression kriging (Hengl et al., 2007) extends ordinary kriging by incorporating auxiliary variables to account for trends in the data that might otherwise violate the stationarity assumption. These auxiliary variables are commonly assumed to have a linear relationship with the variable of interest.
@@
3.4 Dialectometric Analysis
To compare the geographical distribution of speech-to-text model performance against the geographical distribution of dialect similarity, we follow established approaches in dialectometry (Wieling and Nerbonne, 2015) for both distance computation and visualization. Dialectometry (Nerbonne, 2010) is a subfield of linguistics that aims to quantify dialect relations by way of quantitative techniques, which often employs edit distance on word-level data for such a quantification. We compute the aggregate dialect distance pair-wise between sites, where each site has 20 semantically equivalent audio recordings of words, resulting in a site-by-site matrix that is amenable to our visualization method. We present the details below.
3.4.1 Linguistic Distance
For the quantification of linguistic distance between dialect varieties, we follow Bartelds and Wieling (2022) in adopting self-supervised speech representations for the extraction of features on semantically equivalent words, upon which dynamic time warping (DTW) can be applied for a measure of phonetic distance. Formally, let $X$ and $Y$ denote two lists of semantically equivalent words from two dialect varieties, where $X=[w_{X}^{1},w_{X}^{2},\dots,w_{X}^{n}]$ and $Y=[w_{Y}^{1},w_{Y}^{2},\dots,w_{Y}^{n}]$, with $w_{X}^{i}$ and $w_{Y}^{i}$ representing semantically equivalent items. For each pair of items $(w_{X}^{i},w_{Y}^{i})$, we extract their acoustic features using XLSR-53, a self-supervised speech representation model, where we employed an off-the-shelf model finetuned on Italian Common Voice (Ardila et al., 2019) data. [^5] https://huggingface.co/jonatasgrosman/wav2vec2-large-xlsr-53-italian Let $f(w_{X}^{i})$ and $f(w_{Y}^{i})$ represent the feature vectors for the words $w_{X}^{i}$ and $w_{Y}^{i}$, respectively. [^6] We use the last hidden state in our experiments. The acoustic distance $dist(f(w_{X}^{i}),f(w_{Y}^{i}))$ for the $i$-th pair is then computed by way of dynamic time warping, which obtains the distance between two time series that may vary in speed and length by computing the shortest path in a cost matrix. Following Bartelds and Wieling (2022), the distance is normalized by dividing over the length of the shortest path for a fair comparison between sites. The acoustic distance $Dist(X,Y)$ between dialect varieties $X$ and $Y$ is then computed by averaging the pairwise distances across all $n$ items in the lists:
$Dist(X,Y)=\frac{1}{n}\sum_{i=1}^{n}dist(f(w_{X}^{i}),f(w_{Y}^{i}))$
[…]
rowsep=0pt Pearson Spearman chrF2 -0.50 -0.58 BLEU -0.54 -0.50
We apply $Dist(\cdot)$ to all pairwise combinations of dialect sites, resulting in a symmetric site-by-site distance matrix that is amenable to our visualization method described below.
3.4.2 Multidimensional Scaling
Having obtained a site-by-site distance matrix, we employ a visualization method based on dimensionality reduction described in Nerbonne (2010). Nerbonne (2010) shows that a map depicting dialect relations as a continuous surface can be achieved by leveraging classical multidimensional scaling (MDS), [^7] We use the implementation in dialectR (Shim and Nerbonne, 2022). where MDS is a dimensionality reduction method that takes as input a distance matrix and aims to project it to a lower-dimensional space while aiming to preserve the distances in the original high-dimensional space. Formally, given data points $X=\{x_{1},x_{2},\dots,x_{n}\}$, let $D=[d_{ij}]$ be the distance matrix, where $d_{ij}$ represents the distance between $x_{i}$ and $x_{j}$. MDS seeks to minimize the following objective, termed the stress function:
$S(Y)=\sqrt{\frac{\sum_{i,j}(d_{ij}-\|y_{i}-y_{j}\|)^{2}}{\sum_{i,j}d_{ij}^{2}}}$
@@ 4.1 Speech-to-Text Evaluation
rowsep=0pt N Min 1st Q Median 3rd Q Max All Varieties 223 17.97 27.50 35.52 42.51 63.25 Tuscan 8 51.99 57.73 62.32 64.24 72.45 Umbrian 13 48.03 49.67 54.80 60.82 70.45 Abruzzian 8 34.10 38.00 42.89 46.86 53.16 Venetian 31 31.53 37.86 41.76 46.50 57.72 Sicilian 12 34.60 38.17 41.32 43.83 45.56 Ligurian 15 26.88 32.73 37.03 39.77 48.54 Trentinian 11 24.28 31.29 35.52 41.77 48.23 Lucanian 10 23.35 28.73 35.41 42.85 45.12 Molisan 15 25.95 30.94 34.26 39.70 43.26 Apulian 6 21.08 24.30 32.48 38.99 42.20 Friulian 16 22.48 29.40 31.12 37.56 40.81 Ladin 9 17.97 21.72 27.52 28.71 38.10 Piedmontese 13 19.93 25.15 26.40 28.68 30.34 Lombardian 23 18.50 23.50 25.84 30.24 39.07 Sardinian 14 20.04 22.44 23.69 25.13 27.50
@@
rowsep=0pt N Min 1st Q Median 3rd Q Max All Varieties 223 0.00 4.19 7.98 14.50 29.67 Tuscan 8 14.75 20.19 32.72 39.18 49.72 Umbrian 13 12.53 19.45 27.56 34.32 46.08 Abruzzian 8 7.58 12.80 16.14 19.16 27.69 Lucanian 10 3.48 4.31 11.65 14.42 16.75 Sicilian 12 3.45 6.63 11.43 13.70 21.19 Venetian 31 3.31 7.75 10.53 18.28 31.06 Molisan 15 2.20 6.39 8.97 12.50 21.10 Apulian 6 1.53 2.24 8.65 13.14 18.30 Ligurian 15 2.65 4.27 7.45 10.20 17.97 Trentinian 11 3.75 4.38 7.05 10.96 16.51 Friulian 16 1.75 3.02 6.62 11.41 16.96 Lombardian 23 1.62 2.42 4.79 7.43 11.45 Ladin 9 1.36 1.80 4.35 7.04 11.57 Sardinian 14 1.99 2.34 4.30 6.12 8.85 Piedmontese 13 1.74 1.92 3.78 5.90 9.21
@@
Figure 3: Effect of training data size on geostatistical interpolation. Orange is nearest neighbor interpolation (NN); black is inverse distance weighting (IDW); blue is regression kriging (RK).
The chrF results are geographically interpolated and plotted in Figure 2. We observe in the map that zero-shot performance is particularly high in the Tuscan region, where the highest performing dialect site also resides, as shown in Table 4 and Table 5. Prior literature has established standard Italian to be modelled after Tuscan varieties (Hall, 1980; Wieling et al., 2014), which corroborates the trend observed in our results. The similarity of Umbrian dialects with Tuscan dialects is also observable in Umbrian dialects ranking second after Tuscan in Table 4 and Table 5. Furthermore, Table 3 details the correlation between linguistic similarity and the performance scores on the level of dialect sites, where the Pearson correlation for the chrF score is -0.50 and for BLEU -0.54, and the Spearman correlation is -0.58 for chrF and -0.50 for BLEU, suggesting—to answer RQ1—a strong correlation between similarity to the standard (as approximated by the highest performing site) and speech-to-text performance.
@@ 4.3 Geostatistical interpolation
Building on the observation that there is a clear geographical signal in both the performance and dialectometry maps, we next turn to RQ2 and measure to what extent the incorporation of geographical knowledge helps predict zero-shot speech-to-text performance at unseen sites.
[…]
rowsep=0pt NN IDW RK chrF2 9.09 5.86 5.70 BLEU 6.63 4.35 4.19
Table 6: RMSE of geostatistical interpolation methods on unseen sites. NN: Nearest Neighbor, IDW: Inverse Distance Weighting, RK: Regression Kriging.
As shown in Table 6, geostatistical interpolation is highly predictive of both BLEU and chrF scores, with RMSE scores going to as low as 4.19 and 5.70 by regression kriging, our best method. Figure 3 additionally show the effect of training data size on the regression prediction performance. In all results, we observe that the incorporation of distance and covariance between samples as weighting improves over an interpolation by only the value of the nearest neighbor, where regression kriging is consistently the best performing method, followed by inverse distance weighting.
@@ 5.1 Evaluating Dialects as a Continuum
In Figure 2, we observe that for Italian dialects, the evaluation results stand in a continuum that bears similarity to the map of dialect similarity relations. Furthermore, even when one restricts the examined samples to Tuscan varieties, we observe in Table 4 and Table 5 that there is still a disparity that extends the further away one is from the highest performing sites. Our results highlight the need for work in dialect evaluation to take into account the continuous nature of dialects that may exhibit even within dialects often perceived as falling under the same “category”.
While prior work (Kantharuban et al., 2023) has made important headway in individually evaluating non-standard varieties as separate categories and highlighting the performance disparity when compared against the standard, regional variation also exhibits within non-standard varieties such as AAVE (Jones, 2015). Our results suggest that performance prediction in AAVE would arguably also exhibit regional differences, potentially patterning based on how similar the regional varieties are to standard English (e.g. in urban sites), although empirical work is needed to confirm this hypothesis. Importantly, our findings highlight that an evaluation of dialects that is insufficiently balanced geographically therefore carries the risk of overly optimistic views towards model performance at geographically marginal sites, which in turn may lead to social harm towards the subgroups which speak it.
@@ 5.2 Geo-Based Performance Prediction
Extending on the claim of Ahuja et al. (2022), who highlight the role of geography in predicting how well the pivot language for finetuning generalizes to the target language, we explicitly utilize geostatistical methods in our work on geographically proximate dialects. We find both the distance between dialect sites and the covariance between sites to be useful for predicting zero-shot speech-to-text performance at unseen sites. Our results emphasize the geographical structure of dialects, and point to the possibility of leveraging such geographical structure for multilingual transfer between dialects.
@@ 6 Related Work
Kantharuban et al. (2023) stands as the work most similar to our own, where LLMs for both speech and for text are evaluated on the tasks of Machine Translation and Automatic Speech Recognition on regional dialects of high and low-resource languages. They find model performance on such varieties to be highly correlated with lexical and phonetic similarity to the highest performing variety. Their work however considers dialect varieties categorically by considering regional varieties such as Argentinian and Chilean Spanish under the same language of Spanish. Our work instead proposes to work on varieties on a continuous basis across geographically nearby varieties.
With regard to treating varieties as a continuum, Grieve et al. (2024) hypothesize that language models inherently model varieties of language, and propose to leverage sociolinguistic expertise to identify underrepresented varieties. Bafna et al. (2024) model performance degradation of LLMs on closely-related varieties by way of synthetic noise in different linguistic dimensions. In dialectology, the notion of dialects as a continuum is well-established, where computational work in quantifying such a continuum abounds in the field of dialectometry (Heeringa and Nerbonne, 2001; Wieling and Nerbonne, 2015; Nerbonne, 2010).
With regard to bias against dialect varieties in ASR, Feng et al. (2021) show that Dutch ASR systems perform worse on Flemish speakers compared to speakers of all regions in the Netherlands. Kulkarni et al. (2024) show different ASR systems to exhibit different performance biases across 11 states in Brazil, tested on scripted speech and with categorical boundaries between states. Chang et al. (2024) document how self-supervised representations still exhibit a performance disparity in ASR upon AAVE. Our work shows that there is geographical structure in dialect speech-to-text bias that is continuous and correlated with social variables, enabling more fine-grained studies on speech-to-text bias in other languages.
@@ 7 Conclusion
In this paper, we conceptualize dialect relations as a continuum. We find zero-shot performance of speech-to-text systems on dialects to pattern similarly to a measure of similarity to the standard variety, observing strong correlations. We cross-examine our results against established research in linguistic dialectometry. Furthermore, we introduce geostatistical methods that are predictive of zero-shot performance at held-out sites. Our work highlights the need for more research on non-standard varieties that takes into account the continuum nature of dialects. Doing so holds the potential for uncovering bias against speakers of non-standard varieties and helps work toward closing prior evaluation gaps.
@@ 8 Limitations
Our study focuses on related varieties of Italian dialects, which potentially limits the generalizability of our findings. Furthermore, we approximate the standard variety by the best performing dialect variety, which may affect the correlation results depending on how similar the best performing Tuscan variety actually is to standard Italian. Future work should examine how well our insights generalize to other dialect continua, where dialect corpora similar to Vivaldi exist for Sino-Tibetan (Centre for the Protection of Language Resources of China, 2025) and Alpine varieties (Rabanus et al., 2023). Similarly, whether a continuum-based disparity may likewise also exhibit in text-based regional linguistic data remains an open question that future work can explore.
@@ 9 Acknowledgements
We would like to thank the members of the MaiNLP lab and the anonymous reviewers for their invaluable feedback. The icons featured in Figure 2 are from Freepik (Italy, neural network) [^8] https://www.flaticon.com and SVG Repo (pin, arrow, sound wave). [^9] https://www.svgrepo.com Lastly, we recognize the support for SS and BP through the ERC Consolidator Grant 101043235.